OpenStudy (anonymous):

What is the equation of the line passing through (1, 7) and parallel to the line y = –x + 5 in standard form?

OpenStudy (anonymous):

first, we plug the coordinates and slope into the point slope formula: y-7=-1(x-1) simplify: y-7=-x+1 y=-x+8 x+y=8